# ConvTimeNet: A Deep Hierarchical Fully Convolutional Model for Multivariate Time Series Analysis

### v1

@@
Abstract
This paper introduces ConvTimeNet, a novel deep hierarchical fully convolutional network designed to serve as a general-purpose model for time series analysis. The key design of this network is twofold, designed to overcome the limitations of traditional convolutional networks. Firstly, we propose an adaptive segmentation of time series into sub-series level patches, treating these as fundamental modeling units. This setting avoids the sparsity semantics associated with raw point-level time steps. Secondly, we design a fully convolutional block by skillfully integrating deepwise and pointwise convolution operations, following the advanced building block style employed in Transformer encoders. This backbone network allows for the effective capture of both global sequence and cross-variable dependence, as it not only incorporates the advancements of Transformer architecture but also inherits the inherent properties of convolution. Furthermore, multi-scale representations of given time series instances can be learned by controlling the kernel size flexibly. Extensive experiments are conducted on both time series forecasting and classification tasks. The results consistently outperformed strong baselines in most situations in terms of effectiveness. The code is publicly available. [^1] https://github.com/Mingyue-Cheng/ConvTimeNet.
1 Introduction
Time series is a sequence of data points arranged in chronological order. Fully analyzing and mining this type of data Roberts et al. (2013); Kim et al. (2021), such as forecasting evolving trends or classifying the target labels for a given instance, can yield significant gains in various fields, including economics, finance, and biology. Over the past decades, numerous efforts Box et al. (2015); Liu et al. (2021b) have been dedicated to this area. Among these efforts, deep learning-based methods Zhang et al. (2020); Foumani et al. (2023); Liu et al. (2024a) have made significant progress due to their ability to model non-linear relationships and eliminate labor-intensive feature engineering.
@@ 1 Introduction
Over a significant period in the past, the convolutional network He et al. (2016); Zheng et al. (2014); Middlehurst et al. (2023) has played a crucial role in time series analysis, largely due to its inherent properties that strike an excellent balance between computational efficiency and representation quality. Data back to the past years, many representative works Bagnall et al. (2017) of time series analysis typically employ convolutional networks as the backbone. For instance, temporal convolutional network (TCNBai et al. (2018)) and its variants are widely used in modeling temporal variation dependence for the time series forecasting task. Furthermore, a large number of works (such as InceptionTimeIsmail Fawaz et al. (2020), MiniRocketDempster et al. (2021), and MCNNCui et al. (2016)) are also proposed by employing convolutional networks to identify informative patterns from given instances in the classification of time series. However, with the advent of Transformer networks Vaswani et al. (2017); Dosovitskiy et al. (2020), the role of convolutional architecture in time series analysis appears to be diminishing. Recently, Transformer-style networks Wen et al. (2022) have nearly become the dominant role in learning representations of time series, ranging from forecasting Zhou et al. (2021); Liu et al. (2023) to classification tasks Liu et al. (2021a).
In reality, the primary motivation for widely employing the Transformer network is its utilization of the self-attention mechanism. This mechanism enables a more effective capture of long-term sequence dependencies. Despite their effectiveness, using Transformer networks for time series analysis also has some areas of incompatibility. Firstly, multi-scale representations, a crucial element in time series analysis, are challenging to learn within the vanilla Transformer network. This is because the self-attention mechanism can only model fixed-scale representations of sequence input. Secondly, unlike other types of sequence data, both temporal variation Xiao et al. (2022) and cross-variable dependence modeling Fang et al. (2023) are crucial for time series analysis. However, the Transformer network lacks sufficient consideration of modeling interactions across variables, which carry crucial information for accurate predictions of time series. Thirdly, it usually produces quadratic computation complexity Zhou et al. (2021) while attending very long point-wise sequence inputs. This is difficult to avoid because a numerical value at each time step can struggle to reveal informative information, resulting in very long sequence input.
Through careful analysis, we observe that these aforementioned predicaments can largely be avoided by leveraging the inherent properties of convolutional networks. Specifically, firstly, the ability to learn varying scale representations is easily achieved by maintaining a hierarchical convolutional network architecture Cui et al. (2016). Secondly, it is established that the convolutional filter operator naturally complements the modeling of cross-variable relationships Liu et al. (2023). When it comes to temporal dependence, the receptive field can be progressively expanded by stacking deep convolutional operations, thereby acquiring the capacity for global sequence dependence. Thirdly, convolutional neural networks are highly efficient in computation, primarily due to their weight-sharing mechanism Han et al. (2021). This efficiency distinguishes them from Transformer networks, which are constrained by the curse of quadratic time complexity, which is directly proportional to the square of the input size. Along this line, two questions naturally arise: what is the primary obstacle in applying current convolutional networks to time series analysis? If these barriers are overcome, can convolutional networks be revolutionized to achieve superior performance again?
@@
In response to the aforementioned questions, in this work, we propose a novel convolutional network, dubbed ConvTimeNet, designed to serve as a versatile model for time series analysis. One of the most distinctive features of ConvTimeNet is its use of pure convolutional operators to learn the representation of given time series data. Crucially, the network not only preserves the advanced properties of Transformer networks but also inherits several inherent strengths of convolutional networks. Specifically, we first adhere to the modern philosophy of neighboring sequence points together to avoid the sparsity semantic dilemma carried by a single numerical value. Notably, we propose to divide the raw time series into sequences of patches in a data-driven manner, which can better preserve the semantics of local regions. Secondly, on the basis of the building block in Transformer encoder networks, we design a novel fully convolutional block, where deepwise and pointwise convolutions are organized together simultaneously. Particularly, the proposed ConvTimeNet highlights that very deep and hierarchical network architectures are encouraged and very significant in modeling global receptive fields and learning multi-scale representations of given time series instances. To demonstrate the effectiveness and generic capacity of ConvTimeNet, we conduct extensive experiments over a variety of time series analysis tasks including forecasting and classification. The experimental results show that ConvTimeNet could achieve superior or competitive performance compared to strong baselines, including both advanced Transformer networks and pioneering convolutional models. We hope that the proposed ConvTimeNet can serve as an alternative and useful encoder for time series analysis.
2 A Recipe of Fully Convolutional Network for Time Series Analysis
Figure 1: Illustration of the newly proposed ConvTimeNet.
@@ 2 A Recipe of Fully Convolutional Network for Time Series Analysis
2.1 Overall Network Architecture
@@
As illustrated in Figure 1, the forward pass of ConvTimeNet begins with the input of multivariate time series data. Before learning the representation of input data with hidden layers, we develop a deformable patch embedding layer. This layer’s primary role is to transform the raw point-wise time series input into a sequence of patch embeddings. Specifically, the selection of time series points for each patch is determined in a data-driven manner, facilitated by a predictor layer. Subsequently, the transformed time series representation is fed into a stack of full-convolution blocks, with the objective of capturing long-term dependence and learning multi-scale representations. Finally, the architecture culminates in a linear layer, designed to perform time series mining tasks such as forecasting or classification of given time series data.
2.2 Deformable Patch Embedding
@@
The primary objective of time series analysis is to discern the correlation between data points across different time steps. In most existing studies, preserving raw point-wise time series data as input is nearly a default practice. However, a single-time step does not carry as much semantic meaning as a word in a sentence. Consequently, enhancing the information density of each input unit is crucial for improving the final results. We also note that some recent Transformer studies have begun to attempt to extract local information while analyzing their correlations. A common strategy is to aggregate neighboring time points into sub-series level patches. This process, in contrast to modeling each point-wise input individually, actually groups time points into local regions, thereby coupling their dependencies. While this approach is effective, we argue that the fixed-size patch splitting may compromise semantics, leading to reduced performance. For instance, the same pattern in different time series might have varying sizes, a concept supported by well-known shapelet-based approaches Ye and Keogh (2009). Therefore, such strict patch splitting may capture inconsistent information across different instances, potentially leading to sub-optimal results.
Based on the preceding discussion, we assert that time series analysis has experienced success through patch splitting, yet substantial room for improvement remains. To address the dilemma mentioned above, we decided to adopt a philosophy of neighboring time points together, selecting specific time points within each patch in a data-driven manner. This idea of this approach is also consistent with previous work Dai et al. (2017); Chen et al. (2021). As illustrated in the lower left of Figure 1, we additionally introduced a lightweight predictor to adaptively select time points based on given vanilla features, followed by an embedding projection layer. In this study, we refer to this component as the DePatch module.
@@
To achieve this goal, the DePatch module comprises three key steps. Formally, we denote the input time series as $X\in\mathbb{R}^{C\times T}$, where $C$ and $T$ indicate the number of variables (channels) and sequence lengths respectively. Before performing adaptive patch splitting, we first process the raw time series points into $N=\lfloor{\frac{T-P}{S}}\rfloor+2$ fixed-size patches, where the shape of each patch can be denoted by $x\in\mathbb{R}^{P\times C}$. We use the padding operation to ensure the uniform split. For clarity, we omit the superscript ($i$). Here, $P$ denotes the corresponding patch size while $S$ indicates the stride size of the patch window. Subsequently, each patch is assigned three key variables: the center position of this patch $x_{c}$, center offset $\delta_{c}$, and scale variation $\delta_{p}$. It should be noted that $x_{c}$ is decided by the uniform patch splitting above while the $\delta_{c}$ and $\delta_{p}$ are obtained by borrowing the following computation process:
$\delta_{c},\delta_{p}=H(g(x)),$ (1)
@@
$x_{c}^{new}=x_{c}+\delta_{c}$ $,~{}~{}P^{new}=P+2*\delta_{p},$ (2)
$L=x_{c}^{new}-\frac{P^{new}}{2}$ $,~{}~{}R=x_{c}^{new}+\frac{P^{new}}{2},$
@@
where $x_{c}^{new}$ denoted as the new center position of the patch, and $P^{new}$ represents the new length of patch. $L$ and $R$ are denoted as the left and right boundary positions. Finally, based on these two positions, we perform sampling on the original sequence with linear interpolation and obtain a new patch, followed by the projection layer transforms the new patch into the embedding space. With these efforts, the DePatch module allows for the adjustment of each patch’s position and scale based on the input features adaptively, thereby mitigating the semantic destruction caused by hard splitting.
2.3 Fully Convolutional Blocks
@@
To learn high-quality data representation, past methods He et al. (2016); Vaswani et al. (2017) stacked repeated building blocks, a paradigm that is also very popular in the field of data representation learning. Typically, the Transformer block, as a classical design, has been successfully applied in various fields, and these successful applications have inspired methods in the field of time series. However, the Transformer block also has some drawbacks. Firstly, the traditional Transformer block mainly process fixed-scale input, which is not flexible enough. Secondly, the Transformer block is struggling to model the cross-variable dependencies. Finally, the computational complexity of the Transformer block is quadratic related to the length of the input sequence.
@@
Luckily, convolution can effectively address these issues. Motivated by this, we propose a fully convolutional block, as shown in the lower right module of Figure 1. Specifically, we retain the Transformer block’s innovative architecture, incorporating linear and non-linear transformations. The primary distinction lies in our utilization of temporal dependency modeling, achieved through stacking 1D kernel groups, instead of the traditional multi-head attention mechanism. Practically, we employ deepwise convolution as a 1D kernel group, as it offers efficient computational efficiency. Meanwhile we substitute the Feed-Forward Network(FFN) in the Transformer block with a 1x1-sized pointwise convolution to achieve cross-variable dependency modeling. The synergistic combination of these two convolutions effectively completing the comprehensive modeling of the temporal and cross variable dependency.
@@
To achieve a global receptive field, we increase the depth of the model. We discard the traditional residual connection and instead adopt a learnable residual Bachlechner et al. (2021), to circumvent the overfitting issue that often arises from an increased model depth. We also verifiy the effectiveness of the learnable residual mechanism in our experiments. Concurrently, we utilize Batch NormalizationIoffe and Szegedy (2015) in fully convolution block. Furthermore, we employ larger kernels to expand the receptive field. However, in practice, we found that training with large kernels presents a greater challenge. This observation is consistent with recent research findingDing et al. (2022). Therefore, we introduce a re-parameterization mechanism to address this issue. Due to the flexibility of convolutional networks, we can easily achieve a hierarchical architecture by altering the kernel size. We divide the entire deep network architecture into multiple distinct stages, each stage is composed of $K$ fully convolutional blocks. Drawing upon this hierarchical architecture, the time series representation across various scales can be effectively extracted. We also explored the impact of different configurations of hierarchical structures on the model’s capabilities in the experiment.
@@
Concretely, denote $F_{\bm{\theta}}^{l}({\cdot})$ as the larger kernel branch and $G_{\bm{\theta}}^{l}({\cdot})$ as the small kernel branch in $l$-th fully convolution block. Note that $\theta$ represents the weight of the convolution kernel, it will be updated during gradient backpropagation. The re-parameterization mechanism is defined as follows:
${Z}_{DW}^{l}={Gelu}(F_{\bm{\theta}}^{l}({Z^{l-1}})+G_{\bm{\theta}}^{l}({Z^{l-1}})),$ (3)
@@
where $Z^{l-1}$denote the output of $l-1$ convolutional block, $Z_{DW}^{l}\in\mathcal{\mathbb{R}}^{D\times M}$ represents the output deepwise convolution layer in the $l$-th fully convolutional block. $D$ denoted as the number of hidden dimensions and $M$ represents the new sequence length, which is equal to the number of patches $N$. During the inference phase in ConvTimeNet, a sophisticated technique is employed where the weights of the small kernel are zero padded and then combined with those of the larger kernel to form a new, unified convolutional kernel. Denote $\alpha$ as the learnable weights for the residual connections in deepwise convolution layer, typically initialized with 00.
$\hat{Z}_{DW}^{l}=Z^{l-1}+\alpha\times Z_{DW}^{l},$ (4)

### v2

@@
Abstract.
Designing effective models for learning time series representations is foundational for time series analysis. Many previous works have explored time series representation modeling approaches and have made progress in this area. Despite their effectiveness, they lack adaptive perception of local patterns in temporally dependent basic units and fail to capture the multi-scale dependency among these units. Instead of relying on prevalent methods centered around self-attention mechanisms, we propose ConvTimeNet, a hierarchical pure convolutional model designed for time series analysis. ConvTimeNet introduces a deformable patch layer that adaptively perceives local patterns of temporally dependent basic units in a data-driven manner. Based on the extracted local patterns, hierarchical pure convolutional blocks are designed to capture dependency relationships among the representations of basic units at different scales. Moreover, a large kernel mechanism is employed to ensure that convolutional blocks can be deeply stacked, thereby achieving a larger receptive field. In this way, local patterns and their multi-scale dependencies can be effectively modeled within a single model. Extensive experiments comparing a wide range of different types of models demonstrate that pure convolutional models still exhibit strong viability, effectively addressing the aforementioned two challenges and showing superior performance across multiple tasks. The code is available for reproducibility [^1] https://github.com/Mingyue-Cheng/ConvTimeNet.
Time series classification, Deep convolution network
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Time series, sequences of data points arranged in chronological order, are fundamental to various fields due to the valuable insights they provide through analysis and mining (Roberts et al., 2013; Kim et al., 2021). In the realm of computer networks, time series analysis plays a pivotal role, especially amid the rapid expansion of the internet infrastructures. The ability to analyze and interpret temporal patterns in network data is essential for numerous applications. For instance, monitoring network traffic over time aids in detecting anomalies, predicting congestion, and ensuring network security through intrusion detection systems. Moreover, time series analysis enables network administrators to proactively address performance issues by forecasting bandwidth utilization and latency fluctuations, thereby enhancing the efficiency and reliability of network operations.
Over the past decades, numerous efforts (Box et al., 2015; Liu et al., 2021b) have been dedicated to this area. Initially, classical statistical methods ARIMA (Zhang, 2003) were predominant because of their mathematical rigor and interpretability. However, these approaches often struggled with capturing nonlinear patterns and scaling to large datasets. The advent of deep learning introduced data-driven methods such as Long Short-Term Memory (LSTM) networks (Lou et al., 2022) and Gated Recurrent Units (GRUs) (Petneházi, 2019), which modeled sequential data by maintaining hidden states that capture temporal dependencies. Despite advancing the field, these models faced challenges like vanishing gradients and computational inefficiency when handling long sequences.
@@ 1. Introduction
To address these limitations, the convolutional network (He et al., 2016; Zheng et al., 2014; Middlehurst et al., 2023) has been adapted for time series analysis, largely due to its inherent properties that strike an excellent balance between computational efficiency and representation quality. Data from past years shows that many representative works (Bagnall et al., 2017) of time series analysis typically employ convolutional networks as the backbone. For instance, TCN (Bai et al., 2018) and its variants are widely used in modeling temporal variation dependence for the time series forecasting task. Furthermore, a large number of works (such as InceptionTime (Ismail Fawaz et al., 2020), MiniRocket (Dempster et al., 2021), and MCNN (Cui et al., 2016)) are also proposed by employing convolutional networks to identify informative patterns from given instances in the classification of time series. Recently, Transformer-style networks (Wen et al., 2022) have nearly become the dominant role in learning representations of time series, ranging from forecasting (Zhou et al., 2021; Liu et al., 2023) to classification tasks (Liu et al., 2021a). Additionally, novel approaches like Mamba (Gu and Dao, 2023) and adaptations of Graph Neural Networks (GNNs) (Jiang et al., 2023) have emerged, offering alternative perspectives in handling time series data. However, with the advent of Transformer networks (Vaswani et al., 2017; Dosovitskiy et al., 2020), the role of convolutional architecture in time series analysis appears to be diminishing.
Despite the effectiveness of these methods, two fundamental challenges remain unaddressed: local temporal pattern extraction and capturing global sequence dependency. For the formal, the current approach is to uniformly segment the time series into patches, treating each patch as a local pattern. However, this uniform segmentation may not adaptively capture the most informative local patterns, potentially overlooking critical features. Besides, solely focusing on these patterns is insufficient for time series analysis. This is because, in time series classification, local patterns serve as the basic units of representation, but without considering the dependencies and relationships among them on a global scale, it is impossible to distinguish between different time series classes. It is necessary to analyze the dependencies among local patterns from a global perspective. Regarding the latter, although numerous methods have been proposed to model global dependency relationships, they often focus solely on these dependencies while neglecting the effective extraction of local patterns.
@@
In response to the aforementioned questions, in this work, we propose a novel convolutional network, dubbed ConvTimeNet, designed as a versatile model for time series analysis that simultaneously addresses both local pattern extraction and capturing multi-scale repensentation dependency within a single model. One of the most distinctive features of ConvTimeNet is its use of pure convolutional operators to learn the representation of given time series data. Crucially, the network not only preserves the advanced properties of Transformer networks but also inherits several inherent strengths of convolutional networks. Specifically, we first adhere to the modern philosophy of neighboring sequence points together to avoid the sparsity semantic dilemma carried by a single numerical value. Notably, we propose to divide the raw time series into sequences of patches in a data-driven manner, which can better preserve the semantics of local regions. Secondly, we design a novel fully convolutional block where deepwise and pointwise convolutions are organized together simultaneously. Particularly, the proposed ConvTimeNet highlights that very deep and hierarchical network architectures are encouraged and very significant in modeling global receptive fields and learning multi-scale representations of given time series instances. To demonstrate the effectiveness and generic capacity of ConvTimeNet, we conduct extensive experiments over a variety of time series analysis tasks, including forecasting and classification. The experimental results show that ConvTimeNet could achieve superior or competitive performance compared to strong baselines, including both advanced Transformer networks and pioneering convolutional models. We hope that the proposed ConvTimeNet can serve as an alternative and useful model for time series analysis.
Main contributions of this paper can be summarized as follows:
  • We propose ConvTimeNet, a novel fully hierarchical convolutional model that addresses both local pattern extraction and global sequence dependency in time series analysis.
  • We design a convolutional building block that can be deeply stacked by utilizing depthwise and pointwise convolutions along with a reparameterization mechanism.
  • We construct extensive experiments to demonstrate that pure convolutional models can outperform strong baselines on multiple tasks, and numerous additional experiments further confirm the effectiveness of our model design.
2. Related Work
Time series local pattern extraction and capturing its dependency modeling are crucial aspects of time series analysis. Over the years, various approaches have been proposed to model these dependencies, ranging from traditional statistical methods to modern data-driven techniques. Early research predominantly focused on statistical methods such as the ARIMA (Zhang, 2003), which constructs autoregressive models and forecasts using moving averages. ARIMA effectively captures linear relationships within stationary time series data by combining autoregressive terms (AR), differencing operations (I), and moving average terms (MA) to model temporal dependencies. In recent years, deep learning has ushered in a new era of data-driven methods for time series analysis, including those rooted in Convolutional Neural Networks (CNN), Multi-Layer Perceptrons (MLP), and Transformer networks. The CNN-based methods, such as those in (Zheng et al., 2014; Cui et al., 2016), employ sliding convolutional kernels along the temporal dimension to capture temporal dependency. However, these methods have not yielded ideal results in modeling long-range dependencies due to the limited receptive field. On the other hand, MLP-based methods, such as those in (Zeng et al., 2023; Challu et al., 2023), utilize the MLP structure to encode temporal dependencies into the MLP layers. Alternatively, some methods, such as those in (Das et al., 2023), can integrate covariate information into the network.
The Transformer network, renowned for its ability to capture long-range dependencies and cross-variable interactions, is particularly appealing for time series analysis. Consequently, numerous Transformer-based methods have been developed. For instance, Autoformer (Wu et al., 2021) utilizes a self-correlation mechanism to capture temporal dependencies. Crossformer (Zhang and Yan, 2022) is another Transformer-based method that introduces a customized dimension segmentation embedding scheme and an explicit cross-variable attention module designed for forecasting tasks. FormerTime (Cheng et al., 2023a) and GPTH (Liu et al., 2024), on the other hand, employ a hierarchical structure to capture different-scale temporal dependencies and variable dependencies. Meanwhile, the temporal patching operation (Nie et al., 2022) has significantly improved the performance of the model. Despite this, the application of convolutional networks in time series analysis has gradually become less common.
Recently, the application of Large Language Models (LLMs) to time series analysis (Tan et al., 2024; Chang et al., 2023; Cheng et al., 2024) has attracted significant attention in the research community. In the Time-LLM (Jin et al., 2023), the input time series is first tokenized through patching and then aligned with low-dimensional word embeddings using multi-head attention. The aligned outputs, combined with embeddings of descriptive statistical features, are then fed into a frozen, pre-trained language model. The output representations from the language model are flattened and passed through a linear layer to generate the forecast. In this process, the patching employs a uniform segmentation approach.
As modern convolutional techniques continue to evolve, a growing number of convolutional methods are being revisited within the community. For instance, the method (Ding et al., 2022) proposes RepLKNet, which employs a mechanism of re-parameter to circumvent issues associated with expanding the convolutional kernel, thereby addressing the issue of receptive field limitation. Additionally, the TimesNet, introduced recently in (Wu et al., 2022), considers two-dimensional temporal variations generated by periodicities, designed for general tasks in time series analysis.
3. A Recipe of Fully Convolutional Network for Time Series Analysis
Figure 1. Illustration of the newly proposed ConvTimeNet.
[…]
Illustration of the newly proposed ConvTimeNet.
@@ 3. A Recipe of Fully Convolutional Network for Time Series Analysis
3.1. Overall Network Architecture
@@
As illustrated in Figure 1, the forward pass of ConvTimeNet begins with the input of multivariate time series data. To learn the representations of this time series data, we employ a stack of Fully Convolutional Blocks. Leveraging the natural advantages of pure convolution, we vary the sizes of convolutional kernels to deeply model data representations at multiple scales, effectively capturing long-term dependencies. Prior to this, and unlike traditional methods that split patches in a fixed manner, we develop a deformable patch embedding layer. This data-driven patch splitting approach ensures that patterns within each patch are preserved more reasonably, as the selection of time series points for each patch is determined by a predictor layer. Finally, the architecture culminates in a linear layer designed to perform time series mining tasks like forecasting or classification of the given time series data.
3.2. Adaptively Local Pattern Extraction
@@
The primary objective of time series analysis is to discern the correlation between data points across different time steps. In most existing studies, preserving raw point-wise time series data as input is nearly a default practice. However, a single-time step does not carry as much semantic meaning as a word in a sentence. Consequently, enhancing the information density of each input unit is crucial for improving the final results. We also note that some recent Transformer studies have begun to attempt to extract local information while analyzing their correlations. A common strategy is to aggregate neighboring time points into sub-series level patches. This process, in contrast to modeling each point-wise input individually, actually groups time points into local regions, thereby coupling their dependencies. While this approach is effective, we argue that the fixed-size patch splitting may compromise semantics, leading to reduced performance. For instance, the same pattern in different time series might have varying sizes, a concept supported by well-known shapelet-based approaches (Ye and Keogh, 2009). Therefore, strict patch splitting may capture inconsistent information across different instances, likely leading to suboptimal results. To address the dilemma mentioned above, we decided to adopt a philosophy of neighboring time points together, selecting specific time points within each patch in a data-driven manner. This idea of this approach is also consistent with previous work (Dai et al., 2017; Chen et al., 2021). As illustrated in the lower left of Figure 1, we additionally introduced a lightweight predictor to adaptively select time points based on given vanilla features, followed by an embedding projection layer. In this study, we refer to this component as the DePatch module.
@@
To achieve this goal, the DePatch module comprises three key steps. Formally, we denote the input time series as $X\in\mathbb{R}^{C\times T}$, where $C$ and $T$ indicate the number of variables (channels) and sequence lengths, respectively. Before performing adaptive patch splitting, we first process the raw time series points into $N=\lfloor{\frac{T-P}{S}}\rfloor+2$ fixed-size patches, where the shape of each patch can be denoted by $x\in\mathbb{R}^{P\times C}$. Padding is used to ensure a uniform split. For clarity, we omit the superscript ($i$). Here, $P$ denotes the corresponding patch size while $S$ indicates the stride size of the patch window. Subsequently, each patch is assigned three key variables: the center position of this patch $x_{c}$, center offset $\delta_{c}$, and scale variation $\delta_{p}$. It should be noted that $x_{c}$ is decided by the uniform patch splitting above while $\delta_{c}$ and $\delta_{p}$ are obtained by borrowing the following computation process:
(1) $\delta_{c},\delta_{p}=H(g(x)),$
@@
(2) $x_{c}^{new}$ $=x_{c}+\delta_{c},P^{new}=P+2\delta_{p},$
$L$ $=x_{c}^{new}-\frac{P^{new}}{2},R=x_{c}^{new}+\frac{P^{new}}{2}$
@@
where $x_{c}^{new}$ denoted as the new center position of the patch and $P^{new}$ represents the new length of the patch. $L$ and $R$ are denoted as the left and right boundary positions. Based on these two positions, we perform sampling on the original sequence with linear interpolation and obtain a new patch, followed by the projection layer transforms the new patch into the embedding space. With these efforts, the DePatch module allows for the adjustment of each patch’s position and scale based on the input features adaptively, thereby mitigating the semantic destruction caused by hard splitting.
3.3. Multi-Scale Representation Dependency Modeling
@@
To learn high-quality data representation, past methods (He et al., 2016; Vaswani et al., 2017) stacked building blocks, a paradigm that is also very popular in the field of data representation learning. Typically, the Transformer block, as a classical design, has been successfully applied in various fields, and these successful applications have inspired methods in the field of time series. However, the Transformer block also has some drawbacks. On one hand, the traditional Transformer block primarily processes fixed-scale input, which limits its flexibility. On the other hand, the computational complexity of the Transformer block scales quadratically with the length of the input sequence.
@@
We propose a fully convolutional block, as shown in the lower right module of Figure 1. Specifically, we retain the Transformer block’s innovative architecture, incorporating linear and non-linear transformations. The primary distinction lies in our utilization of temporal dependency modeling, achieved through stacking 1D kernel groups instead of the traditional multi-head attention mechanism. Practically, we employ deepwise convolution as a 1D kernel group, as it offers efficient computational efficiency. Meanwhile, we substitute the Feed-Forward Network (FFN) in the Transformer block with a 1x1-sized pointwise convolution to achieve cross-scale dependency modeling. The synergistic combination of these two convolutions effectively completed the comprehensive modeling of the temporal and cross-scale dependency.
@@
To achieve a global receptive field, we increase the depth of the model. We discard the traditional residual connection and instead adopt a learnable residual (Bachlechner et al., 2021), to circumvent the overfitting issue that often arises from an increased model depth. We also verify the effectiveness of the learnable residual mechanism in our experiments. Concurrently, we utilize Batch Normalization (Ioffe and Szegedy, 2015) in a fully convolution block. Furthermore, we employ larger kernels to expand the receptive field. However, in practice, we found that training with large kernels presents a greater challenge. This observation is consistent with recent research finding (Ding et al., 2022). Therefore, we introduce a re-parameterization mechanism to address this issue. We divide the entire deep network architecture into multiple distinct stages, each stage is composed of $K$ fully convolutional blocks. Drawing upon this hierarchical architecture, the time series representation across various scales can be effectively extracted. We also explored the impact of different configurations of hierarchical structures on the model’s capabilities in the experiment.
@@
Concretely, denote $F_{\boldsymbol{\theta}}^{l}(\cdot)$ as the larger kernel branch and $G_{\boldsymbol{\theta}}^{l}(\cdot)$ as the small kernel branch in the $l$-th fully convolutional block. Note that $\boldsymbol{\theta}$ represents the weights of the convolution kernels; they will be updated during gradient backpropagation. The re-parameterization mechanism is defined as follows:
(3) ${Z}_{DW}^{l}={Gelu}(F_{\boldsymbol{\theta}}^{l}({Z^{l-1}})+G_{\boldsymbol{\theta}}^{l}({Z^{l-1}})),$
@@
where $Z^{l-1}$denotes the output of $l-1$ the convolutional block, $Z_{DW}^{l}\in\mathcal{\mathbb{R}}^{D\times M}$ represents the output deepwise convolution layer in the $l$-th fully convolutional block. $D$ denoted as the number of hidden dimensions and $M$ represents the new sequence length, which is equal to the number of patches $N$. During the inference phase in ConvTimeNet, a sophisticated technique is employed where the weights of the small kernel are zero-padded and then combined with those of the larger kernel to form a unified convolutional kernel. Denote $\alpha$ as the learnable weights for the residual connections in the deepwise convolution layer, typically initialized with 00.
(4) $\hat{Z}_{DW}^{l}=Z^{l-1}+\alpha\times Z_{DW}^{l},$